# Probing the pion gluon distribution at small - $x$ in photon-induced interactions at LHC

### v1

@@ II Formalism
In this section, we will present a brief review of the formalism needed to describe the heavy quark photoproduction associated with a leading neutron (HQ+LN) in ultraperipheral collisions. The process is represented in Fig. 1 for $pp$ and $pPb$ collisions, where we also present the main kinematical variables. The basic idea is that in a hadronic collision with center-of-mass energy $\sqrt{s}$, one of the incident protons (or nucleus) can be considered as a source of photons of energy $\omega$, which will interact with the proton target with a squared center-of-mass energy $W^{2}_{\gamma p}=2\omega\sqrt{s}$ upc . Assuming the Sullivan process Sullivan:1971kd , the photon scatters off the pion cloud of the proton target, with the squared center-of-energy for the photon-pion interaction being given by ${W}_{\gamma\pi}^{2}=(1-x_{L})\,W_{\gamma p}^{2}$, where $x_{L}$ is the proton momentum fraction carried by the neutron. Moreover, $t$ is the square of the four-momentum of the exchanged pion, which can be expressed in terms of the measured quantities $x_{L}$ and transverse momentum $p_{T}$ of the neutron, as follows:
@@ III Results
In Fig. 4 we present our predictions for the rapidity distributions for the charm and bottom photoproduction associated with a leading neutron in $pp$ collisions at the LHC, derived assuming different parameterizations for the pion gluon distribution. The results are obtained by calculating Eq. (2). The first term in Eq. (2) is determined by the photon flux for a photon with energy $\omega\propto e^{Y}$ and the heavy quark photoproduction cross-section for a given photon-proton center-of-mass energy $W_{\gamma p}$. While $\sigma_{\gamma p\rightarrow Q\bar{Q}+X+n}$ increases with $W_{\gamma p}$, the photon flux strongly decreases when the photon energy is of the order of $\omega_{\mbox{max}}\approx\gamma_{L}/R_{p}$, becoming almost zero for larger photon energies. As a consequence, this contribution increases with the rapidity up to a maximum and becomes zero at very large $Y$. On the other hand, the second term in Eq. (2) increases for negative values of rapidity, since in this case $\omega\propto e^{-Y}$. For $pp$ collisions, the contributions of both terms in Eq. (2) are identical and symmetric in rapidity. Moreover, we have that the increasing with rapidity is determined by the energy dependence of the heavy quark photoproduction cross-section, being dependent on the pion gluon distribution considered. Therefore, the difference between the predictions presented in Fig.4 is larger for larger values of $|Y|$. However, the results are also distinct at midrapidities ($Y\approx 0$), with the central predictions for the charm production associated with the JAM21 and xFitter parameterizations differing by a factor $\approx 2$. A similar difference is also observed in the case of bottom production (right panel).
@@
In Table 1 we present our results for the charm and bottom photoproduction cross-sections with a leading neutron in $pp$ collisions at $\sqrt{s}=13$ TeV and $pPb$ collisions at $\sqrt{s}=8.1$ TeV, derived assuming different rapidity ranges and using distinct pion PDF parameterizations. In particular, predictions for the rapidity range covered by typical central ($2.0\leq Y\leq 2.0)$ and forward ($2.0\leq Y\leq 4.5)$ detectors are presented. We have that the $pPb$ predictions are $\approx 3$ orders of magnitude larger than for $pp$ collisions. Moreover, the bottom cross-sections are one order of magnitude smaller than for the charm production. Finally, our results indicate that the predictions are sensitive to the pion gluon distribution assumed in the calculations.
@@ Acknowledgements.
V.P.G. and J. T. de Souza were partially supported by CNPq, CAPES (Finance Code 001), FAPERGS and INCT-FNA (Process No. 464898/2014-5).

### v2

@@ II Formalism
In this section, we will present a brief review of the formalism needed to describe the heavy quark photoproduction associated with a leading neutron (HQ+LN) in ultraperipheral collisions. In this exploratory study, we will focus on the direct contribution, in which the photon interacts as a point-like particle, and postpone the inclusion of the resolved processes, where it fluctuates into a hadronic state, for a future publication (For a detailed discussion see, e.g., Ref. Klasen:2002xb ). The direct photoproduction process is represented in Fig. 1 for $pp$ and $pPb$ collisions, where we also present the main kinematical variables. The basic idea is that in a hadronic collision with center-of-mass energy $\sqrt{s}$, one of the incident protons (or nucleus) can be considered as a source of photons of energy $\omega$, which will interact with the proton target with a squared center-of-mass energy $W^{2}_{\gamma p}=2\omega\sqrt{s}$ upc . Assuming the Sullivan process Sullivan:1971kd , the photon scatters off the pion cloud of the proton target, with the squared center-of-energy for the photon-pion interaction being given by ${W}_{\gamma\pi}^{2}=(1-x_{L})\,W_{\gamma p}^{2}$, where $x_{L}$ is the proton momentum fraction carried by the neutron. Moreover, $t$ is the square of the four-momentum of the exchanged pion, which can be expressed in terms of the measured quantities $x_{L}$ and transverse momentum $p_{T}$ of the neutron, as follows:
@@ III Results
In Fig. 4 we present our predictions for the rapidity distributions for the charm and bottom photoproduction associated with a leading neutron in $pp$ collisions at the LHC, derived assuming different parameterizations for the pion gluon distribution. The results are obtained by calculating Eq. (2). The first term in Eq. (2) is determined by the photon flux for a photon with energy $\omega\propto e^{Y}$ and the heavy quark photoproduction cross-section for a given photon-proton center-of-mass energy $W_{\gamma p}$. While $\sigma_{\gamma p\rightarrow Q\bar{Q}+X+n}$ increases with $W_{\gamma p}$, the photon flux strongly decreases when the photon energy is of the order of $\omega_{\mbox{max}}\approx\gamma_{L}/R_{p}$, becoming almost zero for larger photon energies. As a consequence, this contribution increases with the rapidity up to a maximum and becomes zero at very large $Y$. On the other hand, the second term in Eq. (2) increases for negative values of rapidity, since in this case $\omega\propto e^{-Y}$. For $pp$ collisions, the contributions of both terms in Eq. (2) are identical and symmetric in rapidity. Moreover, we have that the increasing with rapidity is determined by the energy dependence of the heavy quark photoproduction cross-section, being dependent on the pion gluon distribution considered. Therefore, the difference between the predictions presented in Fig.4 is larger for larger values of $|Y|$. However, the results are also distinct at midrapidities ($Y\approx 0$), with the central predictions for the charm production associated with the JAM21 and xFitter parameterizations differing by a factor $\approx 2$. A similar difference is also observed in the case of bottom production (right panel). It is important to emphasize that it should be possible to identify the pion-going direction by tagging the forward neutron, which will allow us to use this information to separate the contribution of each term in Eq. (2), allowing a more precise determination of the pion gluon distribution since the ambiguity about what of the incident protons was the photon emitter is eliminated. Such aspect will be explored in a future publication.
@@
In Table 1 we present our results for the charm and bottom photoproduction cross-sections with a leading neutron in $pp$ collisions at $\sqrt{s}=13$ TeV and $pPb$ collisions at $\sqrt{s}=8.1$ TeV, derived assuming different rapidity ranges and using distinct pion PDF parameterizations. In particular, predictions for the rapidity range covered by typical central ($2.0\leq Y\leq 2.0)$ and forward ($2.0\leq Y\leq 4.5)$ detectors are presented. We have that the $pPb$ predictions are $\approx 3$ orders of magnitude larger than for $pp$ collisions. Moreover, the bottom cross-sections are one order of magnitude smaller than for the charm production. Finally, our results indicate that the predictions are sensitive to the pion gluon distribution assumed in the calculations. Such conclusion strongly motivates the improvement of the theoretical description of the process, as well the extension of our analysis for the production of identified heavy-flavour-hadrons, as e.g. $D$- and $B$-mesons.
@@ Acknowledgements.
V.P.G. and J. T. de Souza were partially supported by CNPq, CAPES (Finance Code 001), FAPERGS and INCT-FNA (Process No. 408419/2024-5).